# LLM2CLIP: Powerful Language Model Unlocks Richer Visual Representation

### v4

@@ 3.2 Stage 2: LLM2CLIP Post Fine-tuning
Table 10: LLM2CLIP Training Method Analysis. We experimented with various possibilities for post-training between the LLM and the pretrained CLIP model’s Vision Encoder (CLIP-V) and Text Encoder (CLIP-T). In the figure, ”CL” denotes the contrastive learning loss, and ”Cat” stands for concatenation. Items a–d correspond to the training methods section in Section 3.2. In this experiment, we performed zero-shot image-text retrieval tests on each text encoder, including the new feature projection space formed after concatenation.
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I i2t t2i
Llama-3.1-8B-CC LLM LoRA 17 704 90.5 81.8 61.5 52.4 98.1 98.2 91.0 92.8 85.7 87.4 85.4 82.5
Table 11: Efficiency study of LLM2CLIP. We use two nodes, each equipped with 8 A100 40GB GPUs, and conduct experiments by maximizing GPU memory utilization to achieve the largest possible batch size. We report training hours, batch size, and retrieval accuracy for each experiment. During offloading, we precompute text captions offline into embeddings using the LLM. This approach is a one-time preprocessing step that effectively reduces the experimental overhead.
4 Experiments
4.1 Default Experimental Settings
Default Settings for Caption Contrastive Fine-tuning We choose to use LLaMA 3.1 8B as the LLM integrated into CLIP pre-training, removing the causal attention mask to enable bidirectional attention. No additional adapters are used in caption contrastive finetune, and average pooling is employed to obtain the sentence token embedding. MNTP is not used by default.
Default Settings for LLM2CLIP Cross-modal Post-training After caption contrastive fine-tuning, we replace the original CLIP text encoder with the LLaMA 3.1 8B model to perform cross-modal contrastive learning post-training with the vision encoder. By default, we open the ViT gradients and freeze the LLM gradients, then add an adapter layer composed of four linear layers after the LLM. The design of the Adaptor is the same as FuseMix [35], using a simple inverted bottleneck MLP architecture. By default, we use 15M DreamLIP annotated subsets of CC 3M and CC 12M. For the 60M data setting, we use DreamLIP re-annotated CC 3M and CC 12M, YFCC 15M, and a 30M subset of the LAION dataset. For the 3M setting, we only use the CC 3M subset. We summerize our default experiments setting and implementation details in Appendix A.
4.2 System Comparison
In the system comparison, we used the default settings from  Section 4.1 throughout. However, we also conducted experiments on different pretrained SOTA CLIP models to verify that applying LLM2CLIP post training can significantly enhance their performance.
4.2.1 Zero-Shot English Text-Image Retrieval
Evaluation Dataset. For short-text retrieval, we used the MSCOCO [22] 5K test set and the Flickr [37] 1K test set. For long-text retrieval, we employed datasets from LongCLIP, including a 1K subset of ShareGPT4V [3], the Urban1K [39] dataset, and the DOCCI [26] dataset.
The ShareGPT4V-1M dataset consists of captions generated using GPT-4V and ShareCaptioner, covering images from LAION, CC, SBU [27], and MS COCO. Urban1K includes captions for 1,000 urban scene images, each richly annotated with detailed descriptions. DOCCI contains 1.5K high-resolution images with human-annotated captions and was used for retrieval evaluation. To facilitate observation, we additionally report the average Top-1 image-to-text (I2T) and text-to-image (T2I) accuracy across these five datasets.
LLM2CLIP Improves CLIP for Long- and Short-Text Retrieval. In Table Table 2, we compared OpenAI’s CLIP, EVA02, and Siglip2, and conducted experiments using three types of visual models: ViT-B/16, ViT-L/14, and SoVit/14. Regardless of the model size or resolution shown in the table, LLM2CLIP yields a significant performance boost for both CLIP and EVA02. For example, under EVA02 at 224 resolution, the LLM2CLIP-60M design achieves an average Top-1 retrieval accuracy improvement of +12 and +14.8. Even for SigLIP2—which was pretrained on 40B data—our method still brings an average performance improvement of +9.4 and +10.6; even on the short-caption datasets Flickr and COCO, where SigLIP2 is already very familiar, improvements of +1 and +1.9 are observed. Comparatively, the improvements for long captions are even larger, as the inherent large window of LLMs fully leverages their advantage in understanding long and complex texts.
Training Data Volume Analysis for LLM2CLIP. In Table Table 2, we also conducted experiments with 3M, 15M, and 60M training data volumes on CLIP and EVA02 using ViT-L/14. We found that, whether at 224 or 336 resolution, more data consistently leads to better retrieval performance, for both long and short text retrieval tasks. Furthermore, we observed that the 3M data experiment on EVA02 is already sufficient to yield a noticeable improvement in long text retrieval tasks; however, no clear gain was observed in short text retrieval. We speculate that with a small data volume, the model first needs to compensate for the loss incurred by disrupting the original cross-modal space—a fact evidenced by the significant difference between the 3M and 15M data experiments.
4.2.2 Cross-Lingual Retrieval
LLMs naturally enable strong multilingual capabilities. As in Table 3, We conducted cross-lingual retrieval experiments where CLIP vision encoders were trained exclusively on English text. We successfully endowed EVA02 with multilingual capabilities, which it previously lacked, and significantly enhanced the already robust multilingual performance of SigLip2.
4.2.3 Zero-shot Imagenet Classification and Linear Probe Performance
Evaluation Dataset. We conduct zero-shot retrieval testing on Imagenet, following the CLIP method. We also perform linear probing, fine-tuning the model with the exact same hyperparameters: batch size of 1024, learning rate of 0.1, momentum 0.9, weight decay 0 and total 50 epochs using SGD.
Imagenet Performance Better In Linear Probe. Our method shows a slight drop in performance on Imagenet for zero-shot classification. We hypothesize that this is due to insufficient alignment of the data, resulting in some categories having very low word frequency, as well as the LLM’s limited ability to distinguish short captions. However, in the linear probe experiment, we observe that after supplementing the category information, LLM2CLIP actually leads to a greater improvement. This demonstrates that LLM2CLIP can indeed provide better visual features. One of the current limitations is that scaling the dataset to a larger size might be necessary to resolve the zero-shot classification issue.
4.2.4 Multimodal Large Language Models Performance
Evaluation Dataset. Following LLAVA1.5 [24], we used OpenAI’s CLIP-ViT-L-336 encoder with a simple MLP head to connect to Vicuna-7B. Pretraining included 558K image-caption pairs and 665K visual instruction samples. We fine-tuned the LLAVA1.5 encoder using LLM2CLIP to assess improvements in visual feature extraction.
LLM2CLIP Improves VLLM Performance. As shown in Table 5, LLM2CLIP fine-tuning enhanced multimodal model performance on over 87.5% of benchmarks, with minor losses in only two tasks. This highlights the potential of LLM2CLIP for improving complex image reasoning and understanding tasks.
4.3 Ablation Study: Stage-1 Caption Contrastive Fine-tuning
Experimental Settings. For simplicity, we select Llama 3.1 1B as the default LLM to investigate different design choices for caption contrastive fine-tuning in this section. We use a subset of the CC 3M dataset for caption contrastive fine-tuning. Additionally, we explore augmenting the training data with the Wikitext-103 dataset [25] for MNTP training and the E5 dataset [29] as supplementary pure-text data for caption contrastive fine-tuning.
Structural Design Ablation. We examine various adaptor structures in Table 8. Comparing rows 4 and 5, as well as rows 6 and 7, we observe that neither Linear nor Transformer adaptors significantly impact the performance of subsequent CLIP post-training. Therefore, we default to using no adaptor structure in Stage 1.
Training Method Ablation. We conduct an ablation study of Stage-1 LLM caption contrastive fine-tuning methods in Table Table 9, observing the following key insights: 1. Freezing LLM gradients and relying solely on adaptors yields poor performance, indicating that enabling LoRA is essential. 2. SimCSE is identified as the most critical loss function. MNTP does not significantly enhance SimCSE, and relying solely on MNTP without SimCSE severely deteriorates performance. 3. Supervised SimCSE substantially outperforms unsupervised SimCSE. 4. The performance difference between causal and bidirectional attention mechanisms is negligible.
Ablation for different LLMs. In Table 6, our method achieves significant improvements over these encoders. The result highlighted in red shows that using the original Llama3.1-8B without CC fine-tuning yields notably poor performance, negatively affecting the original CLIP.
4.4 Ablation Study: Stage-2 LLM2CLIP
Adaptor Design. As shown in Table 8, we explore adaptor structures in the Stage-2 LLM2CLIP cross-modal pre-training. Comparing Linear (×$\times 4$) and Transformer (×$\times 1$), their performances are relatively similar, and thus, we select the simpler Linear adaptor structure. Furthermore, in our layer-wise ablation of Linear adaptors, we find that increasing the adaptor size indeed improves performance. Specifically, adaptor performance improves progressively when transitioning from no adaptor to 1, 2 and 4-layer.
Training Method Ablation. We examine several methods of integrating LLMs into CLIP post-training in Table Table 10, with implementation details described in Appendix A. By comparing experiments (a)-(b) and (b)-(c), we find that neither jointly training two losses nor aligning text embeddings significantly improves performance. However, experiment (d), which applies contrastive learning on the concatenated encoder outputs, achieves notable improvements. For simplicity, we do not adopt this design by default but suggest it as a promising direction for future work.
Optimization Strategy Ablation and Efficiency Analysis of LLM2CLIP. As shown in  Table 11, we conducted experiments comparing methods such as LoRA training, frozen LLM online training, and offloading on fixed hardware. Experimental results demonstrate that, with identical computational resources, our proposed offloading method can increase the batch size by tenfold and extend training time by fourfold, ultimately achieving higher performance.
@@ 4.4 Ablation Study: Stage-2 LLM2CLIP
Data Ratio Ablation. In Table Table 7, we experiment with various dataset compositions. The results show a clear trend: higher proportions of dense captions enhance performance in long-text image retrieval tasks, whereas lower dense caption ratios improve short-text retrieval performance. We select a 1:1 ratio by default, as it provides the most balanced improvement for both short and long texts.
@@
5 Conclusion
Given the powerful text understanding capabilities of LLMs, we aim to harness their potential to enhance cross-modal representation, allowing language—the carrier of human knowledge—to play a pivotal role in disseminating the intelligence of LLMs across multiple modalities. This paper successfully proposes an efficient method to seamlessly integrate LLMs into CLIP post-training, demonstrating substantial performance improvements on state-of-the-art multimodal models, including CLIP, EVA02, and SigLip2.
Appendix A Implementation details
Details for Caption Contrastive Fine-tuning
We default to using LoRA for training the LLM, with a rank of 16 and $\alpha=32$. Training is conducted on four nodes, each with 8 NVIDIA A100 40GB GPUs, with an effective batch size of 2048 for one epoch, using a maximum sequence length of 512 tokens for 68 hours. The Adam optimizer is used with a learning rate of $2.5\times 10^{-4}$ and a linear learning rate warm-up for the first 300 steps. We employ bfloat16 quantization, gradient checkpointing, and FlashAttention 2 to optimize GPU memory consumption.
Details for LLM2CLIP Cross-modal Post-training
We employ the AdamW optimizer with a total batch size of 4096 and set the learning rate to $1\times 10^{-5}$ with a cosine decay rate of 0.05.
Details of multi encoder.
When using multiple encoders, we need to align CLIP-V, CLIP-T, and LLM embeddings. We design a one-layer linear projector for each encoder, maintaining the dimensionality at 1024. For the concatenation head scenario, we design a 2048-dimensional head placed after the concatenation of (LLM, CLIP-T) embeddings and another after CLIP-V.
Appendix B Additional results
Training Data Ablation.
In Table Table 12, we explore the impact of different data sources and volumes on Stage-1 LLM caption contrastive fine-tuning. Our findings include: 1. Comparing the combination of 3M Set1 and 1.5M E5 with 4.5M Set1 alone, adding supervised pure-text data from E5 does not provide practical benefits to CLIP’s cross-modal post-training performance and may even slightly degrade performance. 2. For datasets sampled differently from the same distribution (Set1 and Set2), the best results are achieved when Stage 1 and Stage 2 utilize distinct datasets. The overall retrieval performance may depend primarily on the total amount of data seen across both stages; for instance, the performance using Set2 in Stage 1 closely matches that of combining Set1 and Set2.
XM3600 Multilingual Results.
Figure 4 plots radar charts for zero‑shot T2I and I2T retrieval across 36 languages on the XM3600 dataset. Although Siglip2 is already pre‑trained on a large, multilingual corpus, augmenting it with LLM2CLIP—leveraging a large language model to enhance linguistic understanding—still yields a substantial boost in cross‑lingual retrieval performance, even when the fine‑tuning is performed exclusively on English data.

### v5

@@ Appendix D More Analysis
Table A8: LLM2CLIP Training Method Analysis. We experimented with various possibilities for fine-tuning between the LLM and the pretrained CLIP model’s Vision Encoder (CLIP-V) and Text Encoder (CLIP-T). In the figure, ”CL” denotes the contrastive learning loss, and ”Cat” stands for concatenation. Items a–d correspond to the training methods section in Section 3.2. In this experiment, we performed zero-shot image-text retrieval tests on each text encoder, including the new feature projection space formed after concatenation.
Detailed Results for X3600 Multilingual Task.
In Figure 2(a) and Figure 2(b), we visualize the I2T and T2I results for 36 different languages. It can be observed that our method outperforms the original SigLip2 text encoder, despite the latter being trained with 10% of its training samples from 12 billion alt-texts covering 109 languages. In contrast, our method LLM2CLIP was trained exclusively on 15 million English-only samples. This clearly demonstrates the inherent multilingual advantage gained by incorporating an LLM.
Data Ratio Ablation.
@@
In  Table A3, we experiment with various dataset compositions. The results show a clear trend: higher proportions of dense captions enhance performance in long-text image retrieval tasks, whereas lower dense caption ratios improve short-text retrieval performance. We select a 1:1 ratio by default, as it provides the most balanced improvement for both short and long texts.
@@
However, when training exclusively on synthesized long texts, although retrieval performance on long-text tasks can be significantly improved, performance on short-text retrieval tasks tends to decline. We hypothesize two possible reasons for this phenomenon: First, relying entirely on synthesized data might introduce significant biases in the training samples. Second, when the learned feature space is constructed exclusively from long texts, the model may prioritize global semantic matching and overlook fine-grained information such as specific words, potentially negatively affecting short-text retrieval. Therefore, we believe that for contrastive learning tasks focusing solely on long texts, it is important to consider including both global and local levels of text granularity during training like FLIP (Yao et al. 2021), or introducing sufficient hard examples to encourage the model to attend to low-level semantic details.
Optimization Strategy Ablation and Efficiency Analysis of LLM2CLIP.
As shown in  Table A4, we conducted experiments comparing methods such as LoRA training, frozen LLM online training, and offloading on fixed hardware. Experimental results demonstrate that, with identical computational resources, our proposed offloading method can increase the batch size by tenfold and extend training time by fourfold, ultimately achieving higher performance.
Complete Tables and Additional Visualizations.
In Tables Table A5Table A6Table A7, and Table A8, we provide additional detailed results corresponding to Table 6, 7, 8, and 9 in the main text, facilitating a clear analysis of the differences in retrieval performance between long and short texts. The conclusions remain consistent with those presented in the main text. Additionally, Figure A1 illustrates radar plots highlighting the performance improvements achieved by LLM2CLIP on SigLip2 and EVA02.